A solution of salt water raised the boiling point of water from 100°C to 102.5°C. What is the molality of the solution? Assume Kb of water is 0.51°C/m.

1
Expert's answer
2022-04-25T12:27:05-0400

deltaT = Cm x Kb

102.5 - 100 = Cm x 0.51

Cm = 2.5 / 0.51 = 4.9 m